1.849g of MgCl2 is how many moles of MgCl2?
1
Expert's answer
2020-04-15T13:58:34-0400

molar mass of MgCl2 is (24+35.5x2)=24+71=95

moles = "\frac{given \ mass }{molar \ mass }=\frac{1.849}{95}=0.01946"